OpenStudy (anonymous):

c-b=-1 c-a=1 b-a=2 can anyone help to find the value of a,b n c?

OpenStudy (anonymous):

b=2+a from third equation. Plug that b into the b of the first equation. Now you have c's and a's in the first equation. Solve the system with the second equation for a and c. Plug a in to solve for b.

OpenStudy (anonymous):

i already tried,but the c's n a's in the first equation, have the same eq with the second eq , c-a=1.

OpenStudy (anonymous):

that means you can't solve for the system then, so its just b=2+a and c=a+1... check to see if your equations are correct though, seems odd this would be a question on its on